OpenStudy (anonymous):

"Use the second derivative test to to find the points of inflection of f(x)= e^(-x^2). The second derivative is (4x^2 - 2) e^(-x^2)."

OpenStudy (dumbcow):

set 2nd derivative equal to 0 4x^2 -2 = 0 4x^2 = 2 x^2 = 1/2 x = +-sqrt(2)/2 e^-x^2 = 0 -x^2 = ln(0) x has no real solution